OpenStudy (anonymous):

What two angles would I need to add together to get sin(17pi/12) The angles have to be from the pi circle Sum and difference formulas hw Will fan and medal

OpenStudy (anonymous):

The key thing to do is split up 17pi/12 as a sum/difference of two known unit circle values. 17pi/12 = 9pi/12 + 8pi/12 = 3pi/4 + 2pi/3 Now, use the cosine addition identity. cos(a + b) = cos(a)cos(b) - sin(a)sin(b) cos(3pi/4 + 2pi/3) = cos(3pi/4)cos(2pi/3) - sin(3pi/4)sin(2pi/3) = [ -sqrt(2)/2 ] [ -1/2 ] - [ sqrt(2)/2 ] [ sqrt(3)/2 ] = [ sqrt(2)/4 ] - [ sqrt(2)sqrt(3) / 4 ] Merging the radicals in the second fraction, = [ sqrt(2)/4 ] - [ sqrt(6)/4 ] Putting under a common denominator, = [sqrt(2) - sqrt(6)] / 4

OpenStudy (anonymous):

Thank you so much, can you help me with one more, I'll try not take up too much time @Hannah_Waller

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

whats the question? @minisweet4

OpenStudy (anonymous):

sec (-pi/12)

OpenStudy (anonymous):

im not real sure on this one sorry :(

OpenStudy (anonymous):

It's okay, thanks for the help